Cut a board 20 ft long into three pieces so that the longest piece will be three times as long as each of the other two of equal lengths. Find the length of each piece.
Answer by addingup(3677)

Let's call the length of the two equal length pieces x:
x+x+3x= 20
5x= 20
x= 4 Two of the pieces are 4ft long and the third is 4*3=12
Proof:
12+4+4= 20 we have the correct answer